OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3. A. y - 1 = 3(x + 2) B. y - 2 = 3(x + 1) C. y + 2 = 3(x - 1)

OpenStudy (anonymous):

helpp D';

OpenStudy (anonymous):

do you know the point slope equation formula

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

so it would be y+2 = 3 (x-1)

OpenStudy (anonymous):

or C

OpenStudy (anonymous):

you mixed the points up, it should be x in the y spot and y in the x spot

OpenStudy (anonymous):

ohh okay thank you lol

OpenStudy (anonymous):

oh sht no wait it is A im sorry :3

OpenStudy (anonymous):

-.- i told you lmao

OpenStudy (anonymous):

-_____- my brain is fried too.

OpenStudy (anonymous):

lmfao i feel you cx